### v1

Abstract
Reinforcement learning (RL) is a powerful tool for decision-making in uncertain environments, but it often requires large amounts of data to learn an optimal policy. We propose using prior model knowledge to guide the exploration process to speed up this learning process. This model knowledge comes in the form of a model set to which the true transition kernel and reward function belong. We optimize over this model set to obtain upper and lower bounds on the Q-function, which are then used to guide the exploration of the agent. We provide theoretical guarantees on the convergence of the Q-function to the optimal Q-function under the proposed class of exploring policies. Furthermore, we also introduce a data-driven regularized version of the model set optimization problem that ensures the convergence of the class of exploring policies to the optimal policy. Lastly, we show that when the model set has a specific structure, namely the bounded-parameter MDP (BMDP) framework, the regularized model set optimization problem becomes convex and simple to implement. In this setting, we also show that we obtain finite-time convergence to the optimal policy under additional assumptions. We demonstrate the effectiveness of the proposed exploration strategy in a simulation study. The results indicate that the proposed method can significantly speed up the learning process in reinforcement learning.
I INTRODUCTION
Reinforcement learning has shown great success in solving complex decision-making problems in various domains, such as robotics, games, and finance [1]. However, RL often requires large amounts of data to learn an optimal policy, which can be impractical in many real-world applications. This is especially true in environments where data collection is expensive or time-consuming, such as in robotics or healthcare. The need for large datasets comes from the fact that RL algorithms typically explore the environment randomly, causing the agent to visit many suboptimal states and actions [2].
A key opportunity could arise when having access to model knowledge, which is often the case in many scenarios. This information is typically not used in traditional RL algorithms. For instance, we may know that particular states cannot be reached from certain other states or that there is a specific goal state. In this paper, we propose using such available knowledge, giving information on the set of models that the true model belongs to, to guide the exploration of the agent, thus speeding up the learning process.
More specifically, we propose to estimate bounds on the Q-function using optimistic/pessimistic optimization over the available model set. We then use these bounds to guide the agent’s exploration by prioritizing uncertain or promising regions of the state/action space. Additionally, we propose regularizing the model set optimization problem using a database of observed transitions. We show that the Q-function converges to the optimal Q-function under the exploring policy and that the exploring policy converges to the optimal policy in the regularized case. Although these results hold in the general case, we show that for specific model set characterizations, we obtain a practical algorithm with a convex (regularized) model set optimization.
This work is at the intersection of several research areas in reinforcement learning. We will now briefly review the literature on robust reinforcement learning, smart exploration in reinforcement learning, and model-based reinforcement learning, making connections to the content of the present paper.
In robust reinforcement learning, we typically aim to find policies that are robust to model uncertainty. This can be done by optimizing over a set of models, as in [3, 4, 5, 6]. These works typically focus on adversarial optimization over the model set to find a policy that is robust to the worst-case model in the set. Our work differs in that we use the model set to guide the exploration of the agent, rather than to find a robust policy.
Smart exploration in reinforcement learning is a well-studied topic, with many works focusing on how to explore the environment efficiently to speed up the learning process. Some works focus on using uncertainty estimates to guide the exploration, such as [2, 7, 8, 9]. Our work differs in that we use a set of models to guide the exploration, rather than using uncertainty estimates.
Model-based reinforcement learning is a popular approach to speed up the learning process in RL. In model-based RL, the agent learns a model of the environment and uses this model to plan its actions [10, 11]. Our work differs in that we use the model knowledge to guide the exploration of the agent, rather than to plan its actions.
This work also relates to two key ideas in the literature. The first is the bounded-parameter MDP (BMDP) framework [12] (also known as uMDPs or IMDPs) which characterizes the uncertainty in the model set through intervals. The BMDP framework has been used to develop algorithms that are robust to model uncertainty, such as [6, 13, 14]. The present work builds on this idea by using the BMDP framework to guide the exploration of the agent. The second is Bayesian optimization, which is a popular method for optimizing (static) black-box functions [15] by balancing exploration and exploitation in a Bayesian framework. Bayesian RL builds on this by using Bayesian inference to guide the agent’s exploration [16]. Our work uses similar heuristics to guide the exploration of the agent but differs in that we also use model knowledge.
The main contributions of this paper can be summarized as follows.
  • We introduce a novel exploration strategy for reinforcement learning that leverages bounds on the Q-function obtained from a model set. We provide general theoretical guarantees on the convergence of the Q-function to the optimal Q-function under the proposed exploration strategy.
  • We propose a data-regularized version of the model set optimization problem that ensures convergence of the Q-function bounds to the optimal Q-function.
  • We propose a practical algorithm based on the BMDP framework and show that, under specific conditions, we obtain finite-time convergence to the optimal policy.
  • We demonstrate the effectiveness of our approach in a numerical case study.
The remainder of this paper is organized as follows. In Section II, we introduce the problem formulation. In Section III, we present the general framework for the proposed method and detail most of the technical results. Section IV discusses a practical algorithm that makes use of the theoretical results. Section V discusses a brief simulation study and Section VI concludes the paper.
II PROBLEM FORMULATION
We model the decision-making environment as a Markov decision process (MDP), which we define in a general setting to encompass both finite and infinite (continuous) state and action spaces. We assume an infinite-horizon discounted MDP, where the agent interacts with the environment over an infinite number of time steps. This MDP is defined as the tuple $\mathcal{M}=(\mathcal{X},\mathcal{U},P,g,\gamma)$, with the following ingredients:
  • State Space $\mathcal{X}$ is a measurable (Polish) space representing the set of possible states of the environment. We denote the Borel $\sigma$-algebra on $\mathcal{X}$ by $\mathcal{B}(\mathcal{X})$.
  • Action Space $\mathcal{U}$ is a measurable (Polish) space representing the set of possible inputs or actions that the agent can take. We denote the Borel $\sigma$-algebra on $\mathcal{U}$ by $\mathcal{B}(\mathcal{U})$.
  • Transition Kernel $P:\mathcal{X}\times\mathcal{U}\times\mathcal{B}(\mathcal{X})\to[0,1]$ is a Markovian transition kernel, where $P(x,u,\cdot)$ is a probability measure on $\mathcal{X}$ for all $(x,u)\in\mathcal{X}\times\mathcal{U}$. We assume that for every bounded measurable function $h:\mathcal{X}\to\mathbb{R}$, the mapping $(x,u)\mapsto\int_{\mathcal{X}}h(x^{\prime})P(x,u,dx^{\prime})$ is measurable.
  • Reward Function $g:\mathcal{X}\times\mathcal{U}\to\mathbb{R}$ is a bounded reward function, where $g(x,u)$ is the reward obtained by the agent when taking action $u$ in state $x$.
  • Discount Factor $\gamma\in[0,1)$ is the discount factor that determines the importance of future rewards.
The goal in this setting is to find a policy $\pi\in\Pi$, where $\Pi:=\{\pi:\mathcal{X}\to\text{Prob}(\mathcal{U})\mid\pi\text{ is measurable}\}$, such that the expected cumulative reward or value is maximized. The value function $V^{\pi}:\mathcal{X}\to\mathbb{R}$ of a policy $\pi\in\Pi$ is defined as
$V^{\pi}(x)=\mathbb{E}^{\pi}\left[\sum_{k=0}^{\infty}\gamma^{k}g(x_{k},u_{k})\,\bigg{|}\,x_{0}=x\right],$ (1)
where $(x_{k},u_{k})$ denotes the state-action pair at time $k$ under policy $\pi$, and $\mathbb{E}^{\pi}$ denotes the expectation with respect to the probability measure induced by $\pi$ and the transition kernel $P$. A policy’s Q-function $Q^{\pi}:\mathcal{X}\times\mathcal{U}\to\mathbb{R}$ is defined as
$Q^{\pi}(x,u)=g(x,u)+\gamma\int_{\mathcal{X}}V^{\pi}(x^{\prime})P(x,u,dx^{\prime}),$ (2)
where $Q^{\pi}(x,u)$ is the expected cumulative reward of taking action $u$ in state $x$, then following policy $\pi$. We assume that for every $\pi\in\Pi$, the integrals in (1) and in (2) are well-defined. The optimal Q-function $Q^{*}:\mathcal{X}\times\mathcal{U}\to\mathbb{R}$ is defined as
$Q^{*}(x,u)=\sup_{\pi\in\Pi}Q^{\pi}(x,u).$ (3)
An optimal policy $\pi^{*}$ can be obtained from the optimal Q-function as $\pi^{*}(x)\in\arg\max_{u\in\mathcal{U}}Q^{*}(x,u)$. The optimal Q-function can also be defined in terms of the Bellman optimality operator $\mathcal{T}$ as $Q^{*}=\mathcal{T}Q^{*}$, where the Bellman operator $\mathcal{T}:\mathcal{B}(\mathcal{X}\times\mathcal{U})\to\mathcal{B}(\mathcal{X}\times\mathcal{U})$ is defined as
$(\mathcal{T}Q)(x,u)=g(x,u)+\gamma\int_{\mathcal{X}}\sup_{u^{\prime}\in\mathcal{U}}Q(x^{\prime},u^{\prime})P(x,u,dx^{\prime}).$ (4)
In reinforcement learning, our objective is to find the optimal policy $\pi^{*}$ by interacting with the environment and learning from the observed data $\mathcal{D}:={\{}(x_{k},u_{k},x_{k+1}){\}}_{k=0}^{N-1}$. The transition kernel $P$ and the reward function $g$ are assumed to be unknown to the agent. However, we assume that the agent has access to a model set $\hat{\mathcal{M}}$ that the true transition kernel and reward function belong to. This is formalized in the following assumption.
Assumption 1.
The model set $\hat{\mathcal{M}}$ is defined as ^$\hat{\mathcal{M}}=(\mathcal{X},\mathcal{U},\mathcal{P},\mathcal{G},\gamma)$, where $\mathcal{P}$ is a set of transition kernels and $\mathcal{G}$ is a set of reward functions. The true transition kernel $P$ belongs to $\mathcal{P}$, and the true reward function $g$ belongs to $\mathcal{G}$. The model set $\hat{\mathcal{M}}$ is known to the agent.
Thus, the problem is to find the optimal policy $\pi^{*}$ based on the model set $\hat{\mathcal{M}}$ and the observed data $\mathcal{D}$.
III PROPOSED METHOD
We will first discuss a version of the proposed method that fits the general MDP setting introduced in the previous section. This version cannot be directly applied in practice because an optimization over the general model set is intractable. This version presents the theoretical foundation for a more practical version that we will detail in Section IV.
III-A General Framework
Given the model set, we can determine bounds on the Q-function that hold for all models in the set. These bounds are constructed using the Bellman updates, given by
$(\underaccent{\bar}{\mathcal{T}}\underaccent{\bar}{Q})(x,u)$ $=\inf_{g\in\mathcal{G}}g(x,u)+\gamma\inf_{P\in\mathcal{P}}\int_{\mathcal{X}}\underaccent{\bar}{V}(x^{\prime})P(x,u,dx^{\prime}),$ (5)
$(\bar{\mathcal{T}}\bar{Q})(x,u)$ $=\sup_{g\in\mathcal{G}}g(x,u)+\gamma\sup_{P\in\mathcal{P}}\int_{\mathcal{X}}\bar{V}(x^{\prime})P(x,u,dx^{\prime}),$
in which $\underaccent{\bar}{V}$ and $\bar{V}$ are defined as
$\underaccent{\bar}{V}(x):=\sup_{u\in\mathcal{U}}\underaccent{\bar}{Q}(x,u),\quad\bar{V}(x):=\sup_{u\in\mathcal{U}}\bar{Q}(x,u).$ (6)
These relationships and the resulting bounds are related to adversarial MDPs, where the agent aims to minimize the reward in the worst-case scenario, see e.g. [5].
Lemma 2.
Let $\underaccent{\bar}{\mathcal{T}}$ and $\bar{\mathcal{T}}$ denote the Bellman update operators defined in (5). These operators are contraction mappings on $\mathcal{B}(\mathcal{X}\times\mathcal{U})$. Let $\underaccent{\bar}{Q}^{*}$ and $\bar{Q}^{*}$ be the unique fixed points of $\underaccent{\bar}{\mathcal{T}}$ and $\bar{\mathcal{T}}$, respectively, obtained by repeated application of these operators. Then, the fixed points satisfy
$\underaccent{\bar}{Q}^{*}(x,u)\leq Q^{*}(x,u)\leq\bar{Q}^{*}(x,u),\quad\forall(x,u)\in\mathcal{X}\times\mathcal{U},$ (7)
where $Q^{*}$ is the optimal Q-function defined in (3). Consequently, if we define the corresponding value functions as
$\underaccent{\bar}{V}^{*}(x):=\sup_{u\in\mathcal{U}}\underaccent{\bar}{Q}^{*}(x,u)\quad\text{and}\quad\bar{V}^{*}(x):=\sup_{u\in\mathcal{U}}\bar{Q}^{*}(x,u),$ (8)
then we have
$\underaccent{\bar}{V}^{*}(x)\leq V^{*}(x)\leq\bar{V}^{*}(x),\quad\forall x\in\mathcal{X},$ (9)
where $V^{*}(x):=\max_{u\in\mathcal{U}}Q^{*}(x,u)$ is the optimal value function.
Proof.
Since the reward function $g$ is bounded and $\gamma\in[0,1)$, the space $\mathcal{B}(\mathcal{X}\times\mathcal{U})$ of bounded measurable functions, equipped with the supremum norm, is a Banach space. As shown in [17], the standard Bellman operator defined in (4) is a contraction mapping with modulus $\gamma$, i.e.,
$\|\mathcal{T}Q_{1}-\mathcal{T}Q_{2}\|_{\infty}\leq\gamma\|Q_{1}-Q_{2}\|_{\infty},$ (10)
for any bounded functions $Q_{1}$ and $Q_{2}$. Using similar reasoning, the modified operators $\underaccent{\bar}{\mathcal{T}}$ and $\bar{\mathcal{T}}$—which differ from $\mathcal{T}$ only by taking the infimum and supremum over the model set, respectively—are also contractions with the same modulus $\gamma$. Hence, by the Banach fixed-point theorem, each operator has a unique fixed point, namely $\underaccent{\bar}{Q}^{*}$ for $\underaccent{\bar}{\mathcal{T}}$ and $\bar{Q}^{*}$ for $\bar{\mathcal{T}}$.
Since the true MDP $\mathcal{M}$ is contained in the model set $\hat{\mathcal{M}}$ (with $g\in\mathcal{G}$ and $P\in\mathcal{P}$), for any bounded Q-function $Q$ we have
$\underaccent{\bar}{\mathcal{T}}Q(x,u)\leq\mathcal{T}Q(x,u)\leq\bar{\mathcal{T}}Q(x,u),\quad\forall(x,u)\in\mathcal{X}\times\mathcal{U}.$ (11)
The monotonicity of these operators then implies that their unique fixed points are ordered as in (7), where $Q^{*}$ is the fixed point of $\mathcal{T}$, i.e., the optimal Q-function. The value function bounds (9) follow directly from the Q-function bounds (7) and their definition (8).
This completes the proof. ∎
The paper [12] on bounded-parameter MDPs establishes similar guaranteed (optimistic and pessimistic) bounds for the value function in the finite-state case for a particular model set structure. Our result generalizes this idea by considering the Q-function and extending the setting to infinite state spaces. The key is that even in the infinite-state scenario, provided that the boundedness and contraction conditions hold (as detailed in [17]), the same sandwiching property can be proven.
Note that based on the guaranteed Q-bounds, we can in some cases guarantee that certain inputs are suboptimal or optimal. This is formalized in the following proposition.
Proposition 3.
Let $\underaccent{\bar}{Q}^{*}$ and $\bar{Q}^{*}$ be guaranteed lower and upper bounds on the optimal Q-function $Q^{*}$, respectively, such that (7) holds. Then, the input $u$ at state $x$ is guaranteed to be suboptimal if
$\bar{Q}^{*}(x,u)<\underaccent{\bar}{V}^{*}(x),$ (12)
and the input $u$ at state $x$ is guaranteed to be optimal if
$\underaccent{\bar}{Q}^{*}(x,u)\geq\max_{v\in\mathcal{U}\setminus u}\bar{Q}^{*}(x,v).$ (13)
Proof.
For suboptimality, note that $\bar{Q}^{*}(x,u)$ is the best-case return of action $u$, while $\underaccent{\bar}{V}^{*}(x)$ is the best worst-case return across actions at $x$. Hence, if (12) holds, even the most favorable outcome for $u$ is inferior to the least favorable outcome of at least one other action, ensuring that $u$ is suboptimal.
For optimality, if (13) holds, then the worst-case return for $u$ is at least as high as the best-case return for every other action. Since $Q^{*}(x,u)$ lies between these bounds, $u$ must yield the highest possible Q-value at $x$ and is therefore optimal. ∎
In the next section, we will discuss an exploration strategy that leverages the Q-function bounds to guide the agent’s exploration in the environment.
III-B Exploration Strategy
We propose a class of exploring policies to be used during the acquisition of the data that leverages the Q-function bounds. The exploring policy randomly selects the input $u$ at state $x$ according to weighted random sampling with input-dependent weights
$\phi(x,u)=\begin{cases}\xi,&\text{if }\underaccent{\bar}{Q}(x,u)\geq\max_{v\in\mathcal{U}\setminus\{u\}}\bar{Q}(x,v)\\ \beta(x,u),&\text{if }\underaccent{\bar}{Q}(x,u)\neq\bar{Q}(x,u)\text{ and }\bar{Q}(x,u)>\underaccent{\bar}{V}(x)\\ 0,&\text{if }\underaccent{\bar}{Q}(x,u)=\bar{Q}(x,u)\text{ and }\bar{Q}(x,u)\leq\underaccent{\bar}{V}(x)\\ \zeta,&\text{otherwise},\end{cases}$ (14)
with $\phi:\mathcal{X}\times\mathcal{U}\to\mathbb{R}_{\geq 0}$ the sampling weight of input $u$ at state $x$. Here, $\xi\in\mathbb{R}_{>0}$ and $\zeta\in\mathbb{R}_{\geq 0}$ are constants, and $\beta:\mathcal{X}\times\mathcal{U}\to\mathbb{R}_{>0}$. The induced exploration policy $\pi^{\phi}$ is defined by
$\pi^{\phi}(x,u)=\frac{\phi(x,u)}{\int_{\mathcal{U}}\phi(x,v)\,\mu(dv)},$ (15)
provided that $\int_{\mathcal{U}}\phi(x,v)\,\mu(dv)>0$, with $\mu$ a suitable reference measure on $\mathcal{U}$ (e.g., the counting measure if $\mathcal{U}$ is finite).
The first case in (14) implies that the input $u$ at state $x$ is guaranteed to be optimal (by Proposition 3). For this reason, we assign it the positive weight $\xi$. Note that if there exists an input that satisfies this condition, this rules out the possibility that any other inputs satisfy the second condition in (14). The second case implies that the quality of input $u$ at state $x$ is uncertain but that the input has the potential to improve $\underaccent{\bar}{V}$. We assign this choice of input a positive weight $\beta$ that depends on $x$ and $u$. This allows us to potentially assign more weight to the input if it is highly uncertain or has a high Q-value upper bound $\bar{Q}$. Section III-B1 provides heuristic guidance on the choice of $\beta$, inspired by Bayesian optimization. The third case in (14) assigns zero weight to actions with tight Q-bounds that are guaranteed to be suboptimal, while the fourth case assigns a nonnegative weight of $\zeta$ to inputs which are uncertain yet appear to be suboptimal. Note that in many cases, we can select $\zeta=0$ to prune such inputs. However, this pruning might be too aggressive in some cases if we introduce data-driven regularization of the model set optimization, which will be discussed later in this paper. An illustration representing the exploring policy based on the bounds is provided in Fig. 1.
Figure 1: Schematic representation of Q-function bounds. From this particular example, using Proposition 3 we can deduce that the choice $u=3$ is guaranteed to be suboptimal for this state $x$, as its upper bound Q-value $\bar{Q}(x,3)$ is lower than $\underaccent{\bar}{V}(x)$. By (14), the input choice $u=3$ is therefore assigned a weight of $\zeta$. In contrast, actions $u=1$ and $u=2$, for which the bounds do not decisively rule out optimality, receive weight $\beta(x,u)$. Since $\underaccent{\bar}{Q}(x,1)$ is below $\underaccent{\bar}{V}(x)$, the proposed heuristics in Section III-B1 explore the input $u=2$ more frequently.
The convergence properties of standard Q-learning are retained by the proposed exploring policy, as the exploration is based on guaranteed bounds. To show this, we provide the following Theorem.
Theorem 4.
Let $Q$ be the Q-function that is obtained using standard Q-learning, i.e.,
$Q(x_{k},u_{k})\leftarrow(1-\alpha_{k})Q(x_{k},u_{k})+\alpha_{k}\left[r_{k}+\gamma\max_{u^{\prime}\in\mathcal{U}}Q(x_{k+1},u^{\prime})\right],$ (16)
under the data-gathering policy given by the exploring policy $\pi^{\phi}$ defined in (15) with arbitrary parameters $\xi\in\mathbb{R}_{>0},\zeta\in\mathbb{R}_{\geq 0}$ and function $\beta:\mathcal{X}\times\mathcal{U}\to\mathbb{R}_{>0}$. If every state is visited infinitely often (using, e.g., exploring starts) and the learning rate $\alpha_{k}$ satisfies the Robbins-Monro conditions, i.e., $\sum_{k=0}^{\infty}\alpha_{k}=\infty$ and $\sum_{k=0}^{\infty}\alpha_{k}^{2}<\infty$, then the greedy policy with respect to $Q$ converges to the optimal policy with probability 1 as $k\to\infty$
Proof.
Standard Q-learning convergence (e.g., [18, 17]) guarantees that if every state–action pair is updated infinitely often, then $Q$ converges to $Q^{*}$. In our setting, by design, the exploring policy $\pi^{\phi}$ prunes actions that are guaranteed to be suboptimal by (12). Although this means that not every state–action pair is updated infinitely often, it still means that every state $x$, every action $u$ that could be optimal is updated infinitely often (via exploring starts).
Thus, even if $Q(x,u)$ does not converge to $Q^{*}(x,u)$ for all $(x,u)$, the greedy policy $\pi(x)\in\arg\max_{u\in\mathcal{U}}Q(x,u)$ converges to optimal policy $\pi^{*}(x)$ with probability 1, since for every state $x$, the Q-value of every candidate action either converges to its true optimal Q-value or is strictly dominated by another action that does. ∎
In the next section, we detail how to select the weighting function $\beta$ using several heuristics.
III-B1 Choice of Weighting Function $\beta$
Inspired by acquisition functions in Bayesian optimization [15], we can select efficient weighting functions $\beta$. We are interested in improving the best worst-case Q-value $\underaccent{\bar}{V}(x)$ at state $x$. This improvement as a function of the choice of input $u$ is given by
$I(x,u)=\max\left(0,\bar{Q}(x,u)-\underaccent{\bar}{V}(x)\right).$ (17)
By assuming that the probability of Q-values within the bounds $\underaccent{\bar}{Q}$ and $\bar{Q}$ is uniform, we can calculate the probability that a given input $u$ has an improvement greater than 0. This probability of improvement is given by
$\beta(x,u)=$ $~{}\text{Prob}[I(x,u)>0]$ (18)
$=$ $~{}\frac{\max\left(0,\bar{Q}(x,u)-\underaccent{\bar}{V}(x)\right)}{\bar{Q}(x,u)-\underaccent{\bar}{Q}(x,u)}.$
Alternatively, we can take the expected value of the improvement (17) to obtain
$\beta(x,u)=$ $~{}\mathbb{E}[I(x,u)]$ (19)
$=$ $~{}\frac{\left(\max\left(0,\bar{Q}(x,u)-\underaccent{\bar}{V}(x)\right)\right)^{2}}{2\left(\bar{Q}(x,u)-\underaccent{\bar}{Q}(x,u)\right)}.$
These sampling weights prioritize exploring inputs with a large region of their Q-value uncertainty range higher than the current best-worst-case input. Note that we can assume other probability distributions (with bounded support) for the Q-values between $\underaccent{\bar}{Q}$ and $\bar{Q}$. One such example that generalizes the uniform distribution is the Bates distribution, which allows us to assume a higher relative probability concentration around the midpoint of the Q-value bounds.
In the next section, we detail an approach to include data in the method by regularizing the optimization over the model set $\mathcal{P}$ to bias it towards the observed transitions.
III-C Regularization with Observed Data
We propose that by regularizing the optimization over the model set using observed data $\mathcal{D}$, that the bounds converge to the optimal Q-function under mild conditions. We can re-define the Q-bound Bellman updates (5) as
$(\underaccent{\bar}{\mathcal{T}}\underaccent{\bar}{Q})(x,u)$ $=\inf_{g\in\mathcal{G}}g(x,u)+\gamma\int_{\mathcal{X}}\underaccent{\bar}{V}(x^{\prime})\underaccent{\bar}{P}^{*}(x,u,dx^{\prime}),$ (20)
$(\bar{\mathcal{T}}\bar{Q})(x,u)$ $=\sup_{g\in\mathcal{G}}g(x,u)+\gamma\int_{\mathcal{X}}\bar{V}(x^{\prime})\bar{P}^{*}(x,u,dx^{\prime}),$
with
$\underaccent{\bar}{P}^{*}(x,u,dx^{\prime})$ $:=\arg\min_{P\in\mathcal{P}}\int_{\mathcal{X}}\underaccent{\bar}{V}(x^{\prime})P(x,u,dx^{\prime})+\lambda(x,u)d(P,P_{E}),$ (21)
$\bar{P}^{*}(x,u,dx^{\prime})$ $:=\arg\max_{P\in\mathcal{P}}\int_{\mathcal{X}}\bar{V}(x^{\prime})P(x,u,dx^{\prime})-\lambda(x,u)d(P,P_{E}),$
where we assume that the minimum and maximum exist (which is the case if for instance $\mathcal{P}$ is compact). Here, $P_{E}:\mathcal{X}\times\mathcal{U}\times\mathcal{B}(\mathcal{X})\to[0,1]$ denotes the empirical transition kernel obtained from the observed data $\mathcal{D}$, and $\lambda:\mathcal{X}\times\mathcal{U}\to\mathbb{R}_{\geq 0}$ is a regularization parameter. The function $d(P,P_{E})$ is a distance metric between the transition kernel $P$ and the empirical transition kernel $P_{E}$. The distance metric is used to bias the optimized transition kernel towards the observed data, while still being within the model set. The distance metric $d$ should satisfy the following properties:
  • Nonnegativity: $d(P,P_{E})\geq 0$ for all $P$ and $P_{E}$.
  • Zero distance: $d(P,P_{E})=0$ if and only if $P$ is equal to the empirical transition kernel $P_{E}$ obtained from the observed data $\mathcal{D}$.
The variable $\lambda$ determines the trade-off between the model set and the observed data. It should increase with the size of the dataset $\mathcal{D}$ to ensure that the model set is close to the observed data, such that
$\lim_{T(x,u)\to\infty}\lambda(x,u)=\infty,$ (22)
where $T(x,u,x^{\prime})$ the number of observed transitions from state $x$ under action $u$ to state $x^{\prime}$ in the dataset $\mathcal{D}$, and $T(x,u):=\sum_{x^{\prime}}T(x,u,x^{\prime})$.
Additionally, since the rewards are deterministic, the set of reward functions $\mathcal{G}$ can be reduced to a single reward function at a particular state-action pair whenever a reward is observed,
$\mathcal{G}(x_{k},u_{k})=\{r_{k}\}.$ (23)
Together with the regularized model set optimization, we obtain a convergence result which is formalized in the following proposition.
Proposition 5.
Let $\lambda:\mathcal{X}\times\mathcal{U}\to\mathbb{R}_{\geq 0}$ be a regularization parameter satisfying (22), and let $d(P,P_{E})$ be a distance metric such that $d(P,P_{E})\geq 0$ and $d(P,P_{E})=0$ if and only if $P=P_{E}$. Let $\mathcal{P}$ and $\mathcal{G}$ be the set of transition kernels and the set of reward functions, respectively. Assume that the reward function is deterministic, i.e., we can shrink $\mathcal{G}$ through (23) upon observing rewards. If every state-action pair is visited infinitely often, then,
$\lim_{N\to\infty}\underaccent{\bar}{Q}=\lim_{N\to\infty}\bar{Q}=Q^{*}.$ (24)
Proof.
Since every state-action pair is visited infinitely often, $\lambda(x,u)\to\infty$ (by (22)) for every $(x,u)$. The nonnegative metric $d(P,P_{E})$ then forces the regularized optimization in (21) to select $P=P_{E}$, since the regularization term completely dominates the objective. A single visit of every state-action pair is sufficient to ensure that the reward function set $\mathcal{G}=\{g\}$ for every $(x,u)$. By the law of large numbers, $P_{E}$ converges to the true transition kernel $P$, so the regularized update reduces to the standard Bellman update with the true model $(P,g)$. As the standard Bellman operator is a contraction (see, e.g., [17]), it follows that both the lower and upper Q-function bounds converge to $Q^{*}$.
This completes the proof. ∎
Note that for any value of $\lambda$, the Q-bound updates in (5) are still convergent. However, it should be noted that the resulting converged Q-bounds are only guaranteed to bound the optimal Q-function $Q^{*}$ if $\lambda=0$ (by Lemma 2) or if $P_{E}=P$ and $\lambda=\infty$ (by Proposition 5).
Algorithm 1 Q-learning with exploring policy and regularized model set optimization
Calculate Q-function bounds $\underaccent{\bar}{Q}$ and $\bar{Q}$ using (5).
Initialize $Q$ such that $\underaccent{\bar}{Q}\leq Q\leq\bar{Q}$.
for Episode $e=1,2,3,\dots$ do
     Initialize state $x_{0}$.
     for $k=0,1,2,\dots$ until terminal do
         if $\text{rand}()<\epsilon$ then
              Select $u_{k}$ according to exploring policy $\pi^{\phi}$.
         else
              Select $u_{k}=\arg\max_{u\in\mathcal{U}}Q(x_{k},u)$.
         end if
         Apply $u_{k}$ and observe $r_{k}$ and $x_{k+1}$.
         Reduce the reward function set $\mathcal{G}$ such that $g(x_{k},u_{k})=r_{k}$ for all $g\in\mathcal{G}$.
         Update $Q$ using $(x_{k},u_{k},r_{k},x_{k+1})$ and (16).
     end for
     if $e\mod L=0$ then
         Update the Q-bounds using (20) with (21).
         Saturate the Q-function $Q$ to the bounds $\underaccent{\bar}{Q},\bar{Q}$.
     end if
end for
The regularized model set optimization is incorporated into our proposed method, which is summarized in Algorithm 1, where the data-regularized Q-bound iterations are performed every $L$ episodes. To improve learning efficiency, we employ an $\epsilon$-greedy type strategy, balancing exploration/exploitation. Additionally, we saturate the Q-function to the bounds whenever they are updated. By combining our previous results, we obtain the following convergence guarantee for the exploring policy.
Theorem 6.
Let $\pi^{\phi}$ be the exploring policy defined in (15) for which we additionally assume that $\zeta$ is positive. Let $\lambda$ satisfy (22). Suppose that $d(P,P_{E})$ is a distance metric with $d(P,P_{E})\geq 0$ and $d(P,P_{E})=0$ if and only if $P=P_{E}$, and that $\mathcal{P}$ and $\mathcal{G}$ are the model sets for the transition kernel and reward function, respectively. Assume that the reward function is deterministic, i.e., we can shrink $\mathcal{G}$ through (23) upon observing rewards, and that every state is visited infinitely often. Then, under Algorithm 1, the Q-bounds converge to the optimal Q-function and the exploring policy $\pi^{\phi}$ converges to the optimal policy with probability 1 as $k\to\infty$.
Proof.
Almost all the assumptions of Proposition 5 are satisfied. However, the exploring policy $\pi^{\phi}$ in (15) prunes some of the actions, resulting in not every state-action pair being visited infinitely many times. In particular, $\phi$ assigns a positive weight to all actions except actions that are certain and guaranteed to be suboptimal (since $\zeta>0$). Still, since the Q-values of such actions are already certain, this implies that the model set is certain at these state-action pairs. Hence, the exploring policy does not need to explore these actions further, and the regularized model set optimization has the same result regardless of the value of $\lambda$.
For all other state-action pairs, the exploring policy ensures they are explored infinitely often, resulting in $\lambda(x,u)\to\infty$. The regularized model set optimization therefore converges to the true transition kernel for every state-action pair. We obtain (24), similar to Proposition 5. We satisfy (7), which means that the results of Proposition 3 apply. The exploring policy $\pi^{\phi}$ then selects only those actions that are guaranteed to be optimal, implying that it converges to the optimal policy with probability 1.
This completes the proof. ∎
In the next section, we will discuss a practical algorithm derived from the general framework that relates to bounded-parameter MDP models.
IV PRACTICAL ALGORITHM
In this section, we consider the finite state and action space case, where the MDP transitions are reduced to a transition probability tensor $M$, such that
$\text{Prob}(x_{k+1}=x^{\prime}\mid x_{k}=x,u_{k}=u)=M(x,u,x^{\prime}),$ (25)
with $M\in[0,1]^{|\mathcal{X}|\times|\mathcal{U}|\times|\mathcal{X}|}$ and $\sum_{x^{\prime}\in\mathcal{X}}M(x,u,x^{\prime})=1$ for all $(x,u)\in\mathcal{X}\times\mathcal{U}$. The MDP model set $\hat{\mathcal{M}}$ is defined as a bounded-parameter MDP model [12]. The set that the true transition kernel $M$ belongs to is defined as
$\mathcal{P}=\bigg{\{}M~{}\big{|}~{}\underaccent{\bar}{M}\leq M\leq\bar{M},\sum_{x^{\prime}\in\mathcal{X}}M(x,u,x^{\prime})=1,\quad\forall~{}x,u\bigg{\}},$ (26)
for known, well-defined lower and upper bounds $\underaccent{\bar}{M},\bar{M}\in[0,1]^{|\mathcal{X}|\times|\mathcal{U}|\times|\mathcal{X}|}$. The reward function set that the true reward function $g$ belongs to is defined as
$\mathcal{G}=\left\{g~{}\big{|}~{}\underaccent{\bar}{g}\leq g\leq\bar{g},\quad\forall~{}x,u\right\}.$ (27)
In other words, we assume that the true transition kernel and reward function are bounded by known upper and lower bounds.
The distance metric $d$ is chosen to be the Kullback-Leibler divergence between the transition kernel and the observed data, given by
$d(M,M_{E})=\sum_{x^{\prime}\in\mathcal{X}}M(x,u,x^{\prime})\log\left(\frac{M(x,u,x^{\prime})}{M_{E}(x,u,x^{\prime})}\right),$ (28)
where $M_{E}$ is the empirical transition probability tensor from the observed data $\mathcal{D}$ through $M_{E}(x,u,x^{\prime})=\frac{T(x,u,x^{\prime})}{T(x,u)}$.
The resulting Bellman updates are given by
$(\underaccent{\bar}{\mathcal{T}}\underaccent{\bar}{Q})(x,u)$ $=\underaccent{\bar}{g}(x,u)+\gamma\sum_{x^{\prime}\in\mathcal{X}}\underaccent{\bar}{V}(x^{\prime})\underaccent{\bar}{M}^{*}(x,u,x^{\prime}),$ (29)
$(\bar{\mathcal{T}}\bar{Q})(x,u)$ $=\bar{g}(x,u)+\gamma\sum_{x^{\prime}\in\mathcal{X}}\bar{V}(x^{\prime})\bar{M}^{*}(x,u,x^{\prime}),$
with the regularized model optimization given by
$\underaccent{\bar}{M}^{*}(x,u,x^{\prime})$ $=\arg\min_{M\in\mathcal{P}}\sum_{x^{\prime}\in\mathcal{X}}\underaccent{\bar}{V}(x^{\prime})M(x,u,x^{\prime})+\lambda d(M,M_{E}),$ (30)
$\bar{M}^{*}(x,u,x^{\prime})$ $=\arg\max_{M\in\mathcal{P}}\sum_{x^{\prime}\in\mathcal{X}}\bar{V}(x^{\prime})M(x,u,x^{\prime})-\lambda d(M,M_{E}).$
[…]
$\lambda(x,u)=c\sqrt{\frac{T(x,u)}{\log{(|\mathcal{X}||\mathcal{U}|)}/\delta}},$ (31)
where $c$ is a constant, and $\delta$ is a confidence parameter. This formulation is based on Hoeffding’s inequality [19]. This usage is similar to the PAC-MDP framework [7], as implemented, e.g., in the UCRL2 algorithm [20].
The scenario presented in this section is a straightforward extension of the tabular Q-learning algorithm, where we can use the Q-function bounds to guide the agent’s exploration. Note that all the convergence properties of Q-learning and those discussed in the previous sections are retained. In some particular cases, we can even guarantee finite-time convergence of the exploring policy to the optimal policy, as shown in the next section.
IV-A Finite-Time Convergence
We can guarantee finite-time convergence of the proposed exploration strategy under the following conditions:
Assumption 7.
For every $(x,u)\in\mathcal{X}\times\mathcal{U}$, one of the following holds:
  1. 1.
    Deterministic transitions: For all $x^{\prime}\in\mathcal{X}$, $M(x,u,x^{\prime})\in\{0,1\}$.
  2. 2.
    Tight bounds: For all $x^{\prime}\in\mathcal{X}$, $\underaccent{\bar}{M}(x,u,x^{\prime})=\bar{M}(x,u,x^{\prime})$.
Furthermore, $\lambda(x,u)$ is chosen such that it completely relies on the data for the deterministic case, that is, $\lambda(x,u)=\infty$ if 1) holds and $T(x,u)>0$.
The finite-time convergence result is formalized in the following lemma.
Lemma 8.
Let $\pi^{\phi}$ be the exploring policy defined in (15) and let $Q$ be the Q-function from Algorithm 1, using model set $\hat{\mathcal{M}}$ and distance metric $d$ with $d(P,P_{E})\geq 0$ and $d(P,P_{E})=0$ if and only if $P=P_{E}$. Assume that the reward function is deterministic, i.e., we can shrink $\mathcal{G}$ through (23) upon observing rewards. Assume that every state is visited infinitely often (e.g., via exploring starts), and that Assumption 7 holds. Then, the Q-bounds converge to $Q^{*}$ in finite time with probability 1, and consequently the exploring policy $\pi^{\phi}$ converges to the optimal policy in finite time as well.
Proof.
Consider any state-action pair $(x,u)$. The exploring policy ensures that any state-action pair with uncertain Q-bounds is eventually sampled at least once (in finite time). Under Assumption 7:
  1. 1.
    In the deterministic case, a finite number of visits yields the exact empirical kernel $P_{E}(x,u,\cdot)$; with $\lambda(x,u)=\infty$, the optimization in (21) forces $P=P_{E}$, making the Q-bounds exact in finite time.
  2. 2.
    In the tight-bound case, the Q-bound update reduces to the standard Bellman update, since the model set contains only a single model.
Thus, in both cases, the optimistic and pessimistic regularized model set optimization converges to the true transition kernel in finite time. Similarly, the reward function set $\mathcal{G}$ is reduced to the true reward function $g$ in finite time, since we have finite state and action spaces. The Q-bound iterations therefore become equivalent to standard Q-learning updates, since the model uncertainty is completely resolved in finite time. The Q-bounds converge to the optimal Q-function in finite time with probability 1, and using logic similar to the proof of Theorem 6, the exploring policy therefore converges to the optimal policy in finite time with probability 1.
This completes the proof. ∎
Although the general theoretical framework developed in this work applies to both finite and infinite state/action spaces, formulating a practical algorithm for the infinite-state, infinite-action case is considerably more challenging. In practice, one must resort to function approximation techniques to parameterize the Q-function. Such a parametrization typically leads to difficulties when performing optimistic max-max optimization because the overestimation inherent in the optimistic operator can result in divergence issues or unstable learning dynamics [21]. The development of such an algorithm is left for future work.
An effective workaround is to construct a finite abstraction of the original system. By discretizing the state and action spaces, one can apply the proposed exploration strategy and Q-bound updates in a tractable manner, while still capturing the essential dynamics of the underlying continuous problem.
V RESULTS
In this section, we demonstrate the merits of the proposed practical algorithm by applying it to two benchmark reinforcement learning environments, namely the Frozen Lake and Cartpole environments from the OpenAI Gym [22].
V-A Frozen Lake Environment
In the Frozen Lake environment, the agent navigates a gridworld to avoid obstacles and reach a goal state, receiving a reward of 1 upon success and 0 otherwise. The transitions are stochastic due to the slippery nature of the ice, with the agent moving in each perpendicular direction with probability $1/3$.
We define the model set to include knowledge about adjacency in the gridworld. In particular, we assume that all transitions for which the true model has probability zero are known, while all others have probabilities between 0 and 1. The reward function set $\mathcal{G}$ contains only the true reward function $g$, assuming full knowledge of the rewards. This model set definition is similar to what a human player might assume when playing the game, i.e., the agent knows the layout of the gridworld and the locations of obstacles/goals, but not the exact transition probabilities. The results are shown in Fig. 2.
Figure 2: Evaluation performance of the proposed exploration strategy versus standard $\epsilon$-greedy Q-learning in the Frozen Lake environment. Plotted are the median, and the 5th and 95th percentiles of 100 Monte Carlo runs, where the performance is evaluated every 25 episodes by averaging 100 runs of the greedy policy.
We compare standard (randomly exploring) $\epsilon$-greedy Q-learning to two variants of our algorithm: one using the nonregularized exploring policy ($L=\infty$) and one using the regularized version with $L=50$. All experiments use $\xi=1$, $\zeta=0$, $\beta$ as in (19), $c=5$, $\delta=0.05$, $\alpha=0.05$, and $\gamma=0.95$, with $\epsilon$ decaying exponentially from 1 to 0.01. The results indicate that the proposed method converges to the optimal policy much more quickly and consistently, even when the Q-bound iterations are only performed at the beginning. When performing Q-bound iterations every 50 episodes, the proposed method converges even faster, although this comes at an additional computational cost induced by the repeated calculation of Bellman operators.
V-B Cartpole Environment
In the Cartpole environment, the agent is tasked with balancing a pole on a cart by moving the cart left or right, receiving a reward of 1 per time step the pole is balanced. The transitions are deterministic and governed by the physics of the cartpole system. We create finite state and action spaces by discretizing the continuous state and action spaces, resulting in a stochastic transition model. Furthermore, we perform quadratic reward shaping during the training phase to speed up learning.
We define the model set to include knowledge about the cartpole dynamics, i.e., the cartpole is governed by Newtonian physics. In particular, we assume that the dynamics equations are known up to an unknown mass parameter. We obtain the set $\mathcal{P}$ by looping over the set of possible mass values, obtaining the transition probability tensor for each, then taking the max and min over these tensors to obtain $\underaccent{\bar}{M}$ and $\bar{M}$ in (26). The reward function set $\mathcal{G}$ contains only the true reward function $g$, i.e., we assume full knowledge of the rewards. The results are shown in Fig. 3.
Figure 3: Evaluation performance of the proposed exploration strategy versus standard $\epsilon$-greedy Q-learning in the Cartpole environment. Plotted are the median, and the 5th and 95th percentiles of 100 Monte Carlo runs, where the performance is evaluated every 100 episodes by averaging 100 runs of the greedy policy.
We compare the same algorithms ($\epsilon$-greedy, $L=\infty,L=500$) as in Section V-A, with the same $\xi,\zeta,\beta$, and $\delta$. We use $c=100,\alpha=0.03$, $\gamma=0.97$, and $\epsilon$ decaying exponentially from 1 to 0.001. The results are similar to those in the Frozen Lake environment, showing that the regularized method converges fastest and most consistently, followed by the non-regularized version, and then the standard $\epsilon$-greedy Q-learning algorithm. The performance gain due from the regularization is more limited in this environment, which can be attributed to the larger state space dimension.
VI CONCLUSIONS AND FUTURE WORKS
In this paper, we proposed a novel exploration strategy for reinforcement learning that leverages model-based Q-function bounds to guide the agent’s exploration. Our work establishes multiple theoretical results that guarantee the convergence of the proposed exploration strategy to the optimal policy in a general MDP setting. Furthermore, in the finite state and action space case, and under reasonable assumptions on the model set, we obtain a practical algorithm with the same convergence guarantees. We also demonstrated that, under additional assumptions on the transition probabilities, our method achieves finite-time convergence to the optimal policy. The effectiveness of the proposed strategy was validated in several gym environments, where it outperformed standard exploration methods.
For future work, we plan to apply the practical algorithm to a broader range of examples and to extend the theoretical results to derive a practical algorithm for the infinite-state, infinite-action case.